# Collocation methods for nonlinear differential equations on low-rank manifolds

### v2

@@ 6 Numerical examples
and diffusion parameter $\alpha=0.1$. Discretizing $\Omega$ using $n=64$ points in each dimension and approximating derivatives with a Fourier pseudo-spectral method [26], we obtain a semi-discrete version of the Allen-Cahn equation in the form of (2). To assess the accuracy of the low-rank tensor integrators, we computed a reference solution $X(t)$ by integrating the full tensor solution ($64^{3}$ degrees of freedom) to the semi-discrete equation with the four-stage explicit Runge-Kutta (RK4) method using time step-size $\Delta t=10^{-3}$.
@@
We compared the TT-cross integrator presented in Section 5.1 with the step-truncation SVD (ST-SVD) integrator [42] using three different relative truncation tolerances $\delta=10^{-4},10^{-6},10^{-10}$ for determining the solution rank at each time step. We set the solution rank in the TT-cross simulations equal to the ranks obtained from the ST-SVD simulations with truncation tolerances in order to compare the methods for solutions of the same rank. The rank decrease was performed using TT-SVD truncation and the rank increase by sampling more tensor cross indices than singular vectors using the GappyPOD+E algorithm [39] as described in Section 5.3. Time integration for both ST-SVD and TT-cross was performed with Adams-Bashforth 2 and step-size $\Delta t=10^{-3}$.
@@
incurring a cost of $\mathcal{O}(dnr^{6})$ operations at each time $t$. It is possible to accelerate the computation $G(Y,t)$ by carrying out sums and products of TTs with approximate low-rank tensor arithmetic, black-box tensor cross approximation [14], or randomized algorithms [8]. However such algorithms introduce additional errors in the low-rank approximation that can be difficult to control. In comparison, the TT-cross integrator does not require $G(Y,t)$ in a low-rank form and instead evaluates $G(Y,t)$ at $\mathcal{O}(dnr^{2})$ indices. Thus the computational cost of the cubic nonlinearity for TT-cross is negligible compared to the $\mathcal{O}(dnr^{3})$ cost of the TT-cross-DEIM index selection algorithm and evaluating the subtensors of $Y(t)$ required to integrate the system of equations (58). The total CPU-time required to compute the ST-SVD solution with $\delta=10^{-4}$ was 3070 seconds, whereas the TT-cross method required 324 seconds. For $\delta=10^{-6}$, the CPU-time of ST-SVD amounted to 10670 seconds while the TT-cross method took 402 seconds.
@@
where $g(x,y)=\exp(\sin(x)\cos(y))$. Discretizing $\Omega$ using $n=32$ points in each dimension and approximating derivatives with a Fourier pseudo-spectral method [26] we obtain a semi-discrete version of the ADR equation (95) in the form of (2). To assess the accuracy of the low-rank tensor integrators we computed a reference solution $X(t)$ by integrating the full tensor solution ($32^{4}$ degrees of freedom) to the semi-discrete equation with the four-stage explicit RK4 method using time step-size $\Delta t=10^{-3}$.
@@ 7 Conclusions
We introduced new general purpose dynamical low-rank methods for solving nonlinear differential equations on low-rank manifolds. The methods rely on a particular class of oblique projectors onto the tangent space with a low-rank manifold characterized by a cross-interpolation property. Such projectors collocate the differential equation on a low-rank tensor manifold and give rise to efficient time integration schemes that allow us to integrate differential equations defined by vector fields without low-rank structure on low-rank manifolds. To construct the oblique projections we introduced a new index selection algorithm based on the DEIM for constructing well-conditioned interpolatory projectors in the TT format. Furthermore, we showed that such indices also parameterize low-rank TT manifolds and their tangent spaces with cross interpolation. Our numerical results demonstrate that the oblique projections onto the tangent space yield good approximations on the low-rank manifold in the Frobenius norm that are efficiently computed for problems defined by vector fields without low-rank structure. Our proposed methods thus make dynamical low-rank approximation applicable to a broader class of differential equations and facilitate its use in various practical applications.

### v3

@@ 7 Numerical examples
Table 1: CPU-time and accuracy of low-rank methods for integrating the 3D Allen-Cahn equation (96). The ranks were chosen using $\delta=10^{-3}$ and $\delta=10^{-4}$.
Method Average rank $\|\bm{r}(t)\|_{1}$ Runtime (min) Relative Error ($t=10$)
TT-cross AB2 24.2 4.0 $2.5\times 10^{-2}$
ST-SVD AB2 24.2 16.7 $7.13\times 10^{-3}$
i-PS RK4 24.2 23.9 $2.21\times 10^{-2}$
o-PS RK4 24.2 287.6 $7.13\times 10^{-3}$
TT-cross AB2 32.4 4.3 $3.6\times 10^{-3}$
ST-SVD AB2 32.4 27.2 $1.0\times 10^{-3}$
i-PS RK4 32.4 21.8 $3.6\times 10^{-3}$
o-PS RK4 32.4 522.1 $1.0\times 10^{-3}$
@@
and diffusion parameter $\alpha=0.1$. Discretizing $\Omega$ using $n=64$ points in each dimension and approximating derivatives with a Fourier pseudo-spectral method [26], we obtain a semi-discrete version of the Allen-Cahn equation in the form of (2).
@@
We compared the TT-cross integrator presented in Section 6.1 with the step-truncation SVD (ST-SVD) integrator [42] using different relative truncation tolerances $\delta=10^{-3},10^{-4},10^{-6},10^{-10}$ for determining the solution rank at each time step. We set the solution rank in the TT-cross simulations equal to the ranks obtained from the ST-SVD simulations with truncation tolerances in order to compare the methods for solutions computed with the same rank. The rank decrease was performed using TT-SVD truncation and the rank increase by sampling more tensor cross indices than singular vectors using the GappyPOD+E algorithm [39] as described in Section 6.3. Time integration for both ST-SVD and TT-cross was performed with Adams-Bashforth 2 and step-size $\Delta t=10^{-3}$.
@@
incurring a cost of $\mathcal{O}(dnr^{6})$ operations at each time $t$. It is possible to accelerate the computation $G(Y,t)$ by carrying out sums and products of TTs with approximate low-rank tensor arithmetic, black-box tensor cross approximation [14], or randomized algorithms [8]. However such algorithms introduce additional errors in the low-rank approximation that can be difficult to control. In comparison, the TT-cross integrator does not require $G(Y,t)$ in a low-rank form and instead evaluates $G(Y,t)$ at $\mathcal{O}(dnr^{2})$ indices. Thus the computational cost of the cubic nonlinearity for TT-cross is negligible compared to the $\mathcal{O}(dnr^{3})$ cost of the TT-cross-DEIM index selection algorithm and evaluating the subtensors of $Y(t)$ required to integrate the system of equations (65).
@@
In Table 1 we compare the runtime and relative error at time $t=10$ of the low-rank solutions computed with existing methods (ST-SVD AB2, o-PS RK4) with the solutions computed using the proposed methods (TT-cross AB2, i-PS RK4). We consider two different rank-adaptive simulations with ranks determined by $\delta=10^{-3},10^{-4}$ and report the average $1$-norm of the rank vector over all time steps. The TT-cross AB2 method with an average rank of $24.2$ is approximately $4.2$ times faster than the ST-SVD AB2 method at the same rank, while being roughly half an order of magnitude less accurate. With an average rank of $32.4$, the TT-cross AB2 method is approximately $6.3$ times faster than the ST-SVD AB2 method, while being less than half an order of magnitude less accurate. The speedup observed for the projected RK4 methods is even greater, as these methods require more evaluations of the right-hand side, which includes the cubic nonlinearity. The interpolatory RK4 method with an average rank of $24.2$ is approximately $12$ times faster than the o-PS RK4 method at the same rank, while being roughly half an order of magnitude less accurate. With an average rank of $32.4$, the i-PS RK4 method is approximately $24$ times faster than the ST-SVD AB2 method, while being less than half an order of magnitude less accurate.
@@
where $g(x,y)=\exp(\sin(x)\cos(y))$. Discretizing $\Omega$ using $n=32$ points in each dimension and approximating derivatives with a Fourier pseudo-spectral method [26] we obtain a semi-discrete version of the ADR equation (99) in the form of (2).
@@ 8 Conclusions
We introduced new general purpose dynamical low-rank methods for solving nonlinear differential equations on low-rank manifolds. The methods rely on a particular class of oblique projectors onto the tangent space with a low-rank manifold characterized by a cross-interpolation property. Such projectors collocate the differential equation on a low-rank tensor manifold and give rise to efficient time integration schemes that allow us to integrate differential equations defined by vector fields without low-rank structure on low-rank manifolds. To construct the oblique projections we introduced a new index selection algorithm based on the DEIM for constructing interpolatory projectors in the TT format. Furthermore, we showed that such indices also parameterize low-rank TT manifolds and their tangent spaces with cross interpolation. Our numerical results demonstrate that the oblique projections onto the tangent space yield good approximations on the low-rank manifold in the Frobenius norm that are efficiently computed for problems defined by vector fields without low-rank structure. Our proposed methods thus make dynamical low-rank approximation applicable to a broader class of differential equations and facilitate its use in various practical applications.
Declarations of interest
None.